OpenStudy (anonymous):

Let m and n be two positive integers. Show that (36 m+ n)(m+36 n) cannot be a power of 2.

OpenStudy (zzr0ck3r):

what do you get when you multiply them?

OpenStudy (helder_edwin):

\[ \large (36m+n)(m+36)=36(m+n)^2+5^23^27mn=9[4(m+n)^2+5^27mn] \] hence it cannot be a power of 2